OpenStudy (anonymous):

Hello, if somebody knows, how I can build a 5 year coupon bond, with 5 different also COUPON bonds(1y, 2y, 3y, 4y, 5y). What equation should I use?

OpenStudy (anonymous):

That's what I think about this: -B(1,2,3...n) - Bond for 1,2,3...n years. -w - weight in portfolio -C - coupon paid -N - nominal paid I have to construct B(5x) using B(1), B(2), B(3), B(4), B(5) - that's another B(5). So: (N(1)+C(1))*w1+C(2)*w2+C(3)*w3+C(4)*w4+C(5)*w5= C(5x) (N(2)+C(2) )*w2+C(3)*w3+C(4)*w4+C(5)*w5= C(5x) (N(3)+C(3))*w3+C(4)*w4+C(5)*w5= C(5x) (N(4)+C(4))*w4+C(5)*w5=C(5x) (N(5)+C(5))*w5=N5x+C5x